OpenStudy (anonymous):

Solve 4(x + 1) = 4x + 10. 0 6 All real numbers No solution

OpenStudy (anonymous):

Solve -2(2x + 5) - 3 = -3(x - 1). 16 -16 -10 -13

OpenStudy (mayankdevnani):

you posted 2 questions !!

OpenStudy (anonymous):

either srry i have 6 questions left on ws

OpenStudy (anonymous):

its 150 im algebrad out

OpenStudy (triciaal):

the first one has no solution 4x + 4 = 4x + 10 4 is not = 10

OpenStudy (mayankdevnani):

anyway,here is your answer :- \[\large \bf 4x+4=4x+10\]

OpenStudy (mayankdevnani):

\[\large \bf 4x-4x=10-4\] \[\large \bf 10-4=0\] We know that 10-4=6 and it is not equal to zero.

OpenStudy (mayankdevnani):

so answer to your first question is :- No solution !

OpenStudy (mayankdevnani):

understood? @isuckatmath21

OpenStudy (anonymous):

yea its just hard to comprehend when im half asleep

OpenStudy (mayankdevnani):

now for second one :- \[\large \bf -2(2x+5)-3=-3(x-1)\]

OpenStudy (mayankdevnani):

\[\large \bf -4x-10-3=-3x+3\] \[\large \bf -4x-13=-3x+3\]

OpenStudy (triciaal):

@isuckatmath21 you should just go to sleep and get up 10 mins earlier tomorrow. You will benefit more.

OpenStudy (anonymous):

i can only enter the wrokspace 1 time on like everything @triciaal believe my i would if i could

OpenStudy (anonymous):

and thanks again @mayankdevnani

OpenStudy (mayankdevnani):

\[\large \bf -x=16\] \[\large \bf x=-16=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (mayankdevnani):

and if you want to ask more questions,then please take advice :- its totally free of cost hahah i give you answers as well as its solution to 3 question and remaining questions,you can now solve it

OpenStudy (mayankdevnani):

*gave

OpenStudy (anonymous):

mk ill choose wisely

OpenStudy (mayankdevnani):

and for remaining questions,please do efforts on it

OpenStudy (anonymous):

nvm not that one

OpenStudy (mayankdevnani):

good !!

OpenStudy (mayankdevnani):

is this your last question ?

OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x - 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

last question i need help with

OpenStudy (anonymous):

i can do the rest

OpenStudy (mayankdevnani):

what do you think in this question ? @isuckatmath21

OpenStudy (mayankdevnani):

@isuckatmath21

OpenStudy (anonymous):

im thinking

OpenStudy (mayankdevnani):

oh!!

OpenStudy (mayankdevnani):

the answer would be option C

OpenStudy (anonymous):

dang my guess was d

OpenStudy (anonymous):

y

OpenStudy (mayankdevnani):

REMEMBER :- when we open modulus, \[\large \bf \left| x \right|=+x~OR~-x\]

OpenStudy (mayankdevnani):

got it ? @isuckatmath21

OpenStudy (anonymous):

sorta

OpenStudy (mayankdevnani):

what ?

OpenStudy (anonymous):

nothing

OpenStudy (mayankdevnani):

so understood ?

OpenStudy (anonymous):

ya sure goodnight ima finish my last 4 problems and hit the hay

OpenStudy (mayankdevnani):

goooooddd